In a survey of 275 employees of a company regarding their 401(K) investments, the following data were obtained.
145 had investments in stock funds.
91 had investments in bond funds.
75 had investments in money market funds.
49 had investments in stock funds and bond funds.
37 had investments in stock funds and money market funds.
40 had investments in bond funds and money market funds.
72 had investments exclusively in some other vehicle.
(a) How many of the employees surveyed had investments in all three types of funds?

employees
(b) How many of the employees had investments in stock funds only?

employees

Found 2 solutions by Lightning_Fast, ikleyn:
Answer by Lightning_Fast(78)

a.) 0 people invested in all three funds
b.) 145 invested in only stock funds

Answer by ikleyn(53763)

.
In a survey of 275 employees of a company regarding their 401(K) investments, the following data were obtained.
145 had investments in stock funds.
91 had investments in bond funds.
75 had investments in money market funds.
49 had investments in stock funds and bond funds.
37 had investments in stock funds and money market funds.
40 had investments in bond funds and money market funds.
72 had investments exclusively in some other vehicle.

(a) How many of the employees surveyed had investments in all three types of funds?   

     = (275-72) - 145 - 91 - 75 + (49 + 37 + 40) = 18.

     Answer.  18 employees had investments in all three types of funds.


     On how to get this answer, see the lesson 

         - Advanced problems on counting elements in sub-sets of a given finite set

     in this site.



(b) How many of the employees had investments in stock funds only?
 
    = 91 - 49 - 37 + 18 = 23.

   Answer. 23 employees had investments in stock funds only.


    On how to get this answer, see the lesson 

        - Counting elements in sub-sets of a given finite set

    in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Miscellaneous word problems".


Answers by "Lightning_Fast" are incorrect.